OpenStudy (anonymous):

Simplify by combining like terms. 12p + 5m + p - 4m – 6p 7p + m 19p + 9m 6p + 9m 7p – m

OpenStudy (hoblos):

12p + 5m + p - 4m – 6p = p(12+1-6) +m(5-4) = 7p +m